OpenStudy (trisaba):

How do you get 23 25 27 and 29 in 4 4s no more no less

OpenStudy (trisaba):

im new

OpenStudy (trisaba):

i will give medals

OpenStudy (trisaba):

@dan815

OpenStudy (dan815):

what do yo mean in 4 4s?

OpenStudy (trisaba):

four fours like the way to get 16 is \[4+4+4+4\]

OpenStudy (dan815):

ah okay so divide each of those numbers by 5

OpenStudy (dan815):

by 4*

OpenStudy (dan815):

23/4 you can rewrite 23 as 20+3 (20+3)/4 =20/4 + 3/4 = 5+3/4 so each of the part will be 5 and 3/4th

OpenStudy (trisaba):

?

OpenStudy (dan815):

it doesnt make sense?

OpenStudy (dan815):

okay for example you want to split 16 into 4 4s right

OpenStudy (trisaba):

yeah

OpenStudy (dan815):